A girl of mass m_1 = 60.0 kilograms springs from a
Answer #1

A)
use conservation of energy ,
initial kinetic energy = final kinetic energy + final potential energy
0.5*m*vi^2 = m*g*h + 0.5*m*vbefore^2
0.5*vi^2 = g*h + 0.5*vbefore^2
0.5*8^2 = 9.8*2 + 0.5*vbefore^2
vbefore = 4.98 m/s

B)
use conservation of momentum,
mbefore*vbefore = m after * vafter
60*4.98 = (60+15)*vafter
vafter = 3.98 m/s